# Spend Wisely: Maximizing Post-Training Gains in Iterative Synthetic Data Boostrapping

### v1

@@ B.4 Proofs
with probability at least $1-\sum_{t=0}^{T^{*}(\pi_{\mathrm{exp}},\varepsilon)-1}n_{t}^{-4}$. Then,
$C(\pi_{\mathrm{exp}},T^{*}(\pi_{\mathrm{exp}},\varepsilon))$ $\leq C(\pi_{\mathrm{exp}},T_{1})$
$=\sum_{t=0}^{T_{1}-1}(c_{\mathrm{g}}r_{t}^{-1}+c_{\mathrm{t}})\left(2\left(16\beta-2c\right)c^{-1}C_{L}C_{\Theta}\log c_{\Theta}\right)^{2}\left(\tilde{\mathbb{P}}l_{\theta}^{(0)}\right)^{-2}\left(1-\frac{\gamma c}{8\beta}\right)^{-2t}$
$\leq(c_{\mathrm{g}}r_{0}^{-1}+c_{\mathrm{t}})\left(2\left(16\beta-2c\right)c^{-1}C_{L}C_{\Theta}\log c_{\Theta}\right)^{2}\left(\tilde{\mathbb{P}}l_{\theta}^{(0)}\right)^{-2}\sum_{t=0}^{T_{1}-1}\left(1-\frac{\gamma c}{8\beta}\right)^{-2t}$
$\leq(c_{\mathrm{g}}r_{0}^{-1}+c_{\mathrm{t}})\left(2\left(16\beta-2c\right)c^{-1}C_{L}C_{\Theta}\log c_{\Theta}\right)^{2}\left(\tilde{\mathbb{P}}l_{\theta}^{(0)}\right)^{-2}\frac{(1+\zeta)^{2(T_{1}-1)-1}}{2\zeta+\zeta^{2}}$
$\leq C_{\mathrm{exp}}\varepsilon^{-2}-c_{\mathrm{exp}}.$
for some positive constants $C_{\mathrm{exp}}$ and $c_{\mathrm{exp}}$ that do not depend on $\varepsilon$. So we have proven the last inequality in Equation 9.
Next, we consider the second inequality in Equation 9. When $\pi^{\prime}$ is a constant policy, Theorem B.4 implies that for sufficiently small $\varepsilon$, $T^{*}(\pi^{\prime},\varepsilon)=\infty$ and the inequality holds vacuously. Otherwise, for a polynomial growth policy $\pi^{\prime}$ under assumptions B.1, B.2 and B.3, there exist $\Theta$, $\Theta^{*}$, $l(x;\theta)$, and $\mathcal{R}(x)$ such that for $\theta^{(t)}_{\pi^{\prime}}$,
~$\tilde{\mathbb{P}}l_{\theta^{(t+1)}_{\pi^{\prime}}}-\tilde{\mathbb{P}}l_{\theta^{(t)}_{\pi^{\prime}}}=\min\left\{-\frac{\gamma c}{4\beta}\tilde{\mathbb{P}}l_{\theta^{(t)}_{\pi^{\prime}}}+\rho\gamma\left(2-\frac{c}{4\beta}\right)2C_{L}C_{\Theta}\log c_{\Theta}\frac{1}{\sqrt{n_{t}^{\prime}}},0\right\},$
for some $0<\rho\leq 1$, and by assumption B.3,
@@
Because
~$\tilde{\mathbb{P}}l_{\theta^{(T^{*}(\pi^{\prime},\varepsilon))}_{\pi^{\prime}}}\leq c_{3}^{-1}\left(r^{*}-\mathbb{E}_{x\sim\mathcal{P}_{\theta^{(T^{*}(\pi^{\prime},\varepsilon))}_{\pi^{\prime}}}}\mathcal{R}(x)\right)\leq c_{3}^{-1}\varepsilon,$
@@
so the second inequality in Equation 9 also holds.
@@
Appendix C Algorithm for Realistic Experiments
We follow the algorithm in Algorithm 1. Empirically, in the two realistic settings (image denoising and math reasoning), our iterative learning policy satisfies that $n_{t}$ is an integer multiple of the batch size. Specifically,
@@ Appendix C Algorithm for Realistic Experiments
Appendix D Diffusion Probabilistic Models
To model the reverse process, the time interval $[0,1]$ is discretized into steps $10=\tau_{0}<\tau_{1}<\cdots<\tau_{K}=$ 1. Starting at $\tau_{K}$, the reverse diffusion process gradually removes noise step by step until reaching $\tau_{0}$. At each step $k$, the noise added during the forward process is estimated and removed using a trained noise prediction model $\bm{\varepsilon}_{\theta}(\cdot,\tau_{k})$. This model learns to predict the noise $\bm{\varepsilon}$ based on the noisy sample $\bm{x}_{\tau_{k}}$, enabling the reconstruction of cleaner samples at each step.
where $\mathscr{D}$ denotes the distribution of training data.
@@ E.1 Toy Example
In this subsection, we extend the theoretical intuition from Theorem 3.1 to high-dimensional settings and then derive the expected reward of each iteration, as well as the corresponding optimal expected reward.
@@
Under the MLE updation, suppose that the generator is a $d$-dimensional Gaussian distribution $\mathcal{N}(\bm{\theta},\sigma^{2}\bm{I}_{d})$, the reward model is an exponential function $\mathcal{R}(\bm{x})=\exp\left(-\lVert\bm{x}\rVert^{2}/(2\kappa^{2})\right)$, and $\left\|\bm{\theta}^{(0)}\right\|\leq(1+\sigma^{2}/\kappa^{2})^{T}(d(\sigma^{2}+\kappa^{2}))^{1/2}$. Considering the optimization problem
$\max_{\pi}$ $\quad\mathbb{E}_{\bm{\theta}^{(T)}}[\mathbb{E}_{\bm{x}\sim\mathcal{P}_{\bm{\theta}^{(T)}}}\mathcal{R}(\bm{x})]$
s.t. $\quad\sum_{k=0}^{T-1}n_{t}\leq C,$
@@
Since each component of $\bm{\theta}^{(t)}$ is i.i.d. and its distribution conditioned on the initial parameter $\theta^{(0)}$ and the policy $\pi$ is given by Lemma A.1, it follows that
$p(\bm{\theta}^{(T)}\mid\bm{\theta}^{(0)},\pi)=\mathcal{N}\left(\frac{\bm{\theta}^{(0)}}{(1+\sigma^{2}/\kappa^{2})^{T}},\sigma^{2}\sum_{t=0}^{T-1}\frac{1}{n_{t}(1+\sigma^{2}/\kappa^{2})^{2(T-t)-1}}\bm{I}_{d}\right)$
for a specific policy $\pi:\{n_{t}\}_{t}$. By the definition of the reward model, $\mathcal{R}(x)$ factors as $\mathcal{R}(\bm{x})=\prod_{i=1}^{d}\mathcal{R}(x_{i})$ with $\bm{x}=(x_{1},x_{2},\cdots,x_{d})$. Repeating the argument in the proof of Theorem 3.1 shows that
@@
$=\left(\frac{\kappa}{\sqrt{\sigma^{2}+\kappa^{2}+\sigma_{T}^{2}}}\right)^{d}\exp\left(-\frac{\lVert\bm{\theta}^{(0)}\rVert^{2}}{2(\sigma^{2}+\kappa^{2}+\sigma_{T}^{2})(1+\sigma^{2}/\kappa^{2})^{2T}}\right).$
With the assumption $\lVert\bm{\mu}_{T}\rVert\leq\left(d(\sigma^{2}+\kappa^{2})\right)^{1/2}$, the right-hand side increases as $(\sigma^{2}+\kappa^{2}+\sigma_{T}^{2})^{-1/2}$ increases. Hence, figuring with the proof of Theorem 3.1 in Appendix A, one obtains
@@
Theorem E.1 therefore shows that the optimal iterative policy in the high-dimensional case mirrors the one-dimensional scenario. Next, we derive the expected reward in the high-dimensional case. Specifically, if $\bm{x}\sim\mathcal{N}(\bm{\theta},\sigma^{2}\bm{I}_{d})$, then
$\mathbb{E}_{x\sim\mathcal{N}(\bm{\theta},\sigma^{2}\bm{I}_{d})}[\mathcal{R}(\bm{x})]$ $=\int\frac{1}{(2\pi\sigma^{2})^{d/2}}\exp\left(-\frac{\lVert\bm{x}\rVert^{2}}{2\kappa^{2}}-\frac{\lVert\bm{x}-\bm{\theta}\rVert^{2}}{2\sigma^{2}}\right)\mathrm{d}x$
@@
$L(\bm{\theta}^{(t)})=|S|^{-1}\sum_{\bm{x}\in S}\left\|\bm{\varepsilon}_{\bm{\theta}^{(t)}}(\alpha_{\tau_{\bm{x}}}\bm{x}+\beta_{\tau_{\bm{x}}}\bm{\varepsilon}_{\bm{x}},\tau_{\bm{x}})-\bm{\varepsilon}_{\bm{x}}\right\|^{2}$
@@
E.2 Image Denoising
We fixed the batch size $B=640$ and the learning rate as $5\times 10^{-5}$ for all the experimental settings. For each iteration, the synthetic data are continuously generated, selected by the reward model $\mathcal{R}$, and then added into a chunk until the chunk size reaches $n_{t}$. The detailed procedure for our training is listed in Algorithm 2. All the codes for training and plotting for the image denoising experiments are available on GitHub: https://github.com/zylipku/IterativeImaging.
To demonstrate the advantages of increasing policies, particularly exponential growth policies, we evaluate computational costs in terms of floating-point operations (FLOPs), which are directly proportional to the number of network inferences. In diffusion models, generating a single sample requires multiple forward passes, and each optimization step incurs an additional forward pass.
E.3 Math Reasoning
@@
6:        Sample a mini-batch =$\mathcal{B}=\{(q_{j},a_{j})\}_{j}\subseteq\mathcal{D}$.
@@
$L(\theta)=\frac{1}{|\mathcal{B}|}\sum_{(\hat{a},q)\in\mathcal{B}}\log p(\hat{a}\mid q;f(\cdot,\theta)).$
@@
We summarize the algorithm for iterative learning on math reasoning in Algorithm 3.
The hyperparameters are specified as follows. For each iteration, we only train 1 epoch with the selected data. We set the learning rate as $10^{-7}$ with the constant learning rate scheduler. The training batch size $B=256$, and we totally update the generator for around 1000 training steps. We set the temperature as 0.3 during generation to provide diversity of answers, while set it as 0 during evaluation to ensure accuracy. We set the maximum generation length as 512. The few-shot examples are followed with the ICL generation setting of GSM8k dataset in OpenCompass (Contributors, 2023). We develop our code based on the publicly available OpenRLHF framework (Hu et al., 2024). All the codes for training and plotting for the math reasoning experiments are available on GitHub: https://github.com/yangpuPKU/IterativeLearning.

### v2

@@ B.4 Proofs
with probability at least $\left(1-\sum_{t=0}^{T^{*}(\pi_{\mathrm{exp}},\varepsilon)-1}n_{t}^{-4}\right)$. Then the equality
$\begin{aligned} C(\pi_{\mathrm{exp}},T^{*}(\pi_{\mathrm{exp}},\varepsilon))&\leq C(\pi_{\mathrm{exp}},T_{1})\\ &=\sum_{t=0}^{T_{1}-1}(c_{\mathrm{g}}r_{t}^{-1}+c_{\mathrm{t}})\left(2\left(16\beta-2c\right)c^{-1}C_{L}C_{\Theta}\log c_{\Theta}\right)^{2}\left(\tilde{\mathbb{P}}l_{\theta}^{(0)}\right)^{-2}\left(1-\frac{\gamma c}{8\beta}\right)^{-2t}\\ &\leq(c_{\mathrm{g}}r_{0}^{-1}+c_{\mathrm{t}})\left(2\left(16\beta-2c\right)c^{-1}C_{L}C_{\Theta}\log c_{\Theta}\right)^{2}\left(\tilde{\mathbb{P}}l_{\theta}^{(0)}\right)^{-2}\sum_{t=0}^{T_{1}-1}\left(1-\frac{\gamma c}{8\beta}\right)^{-2t}\\ &\leq(c_{\mathrm{g}}r_{0}^{-1}+c_{\mathrm{t}})\left(2\left(16\beta-2c\right)c^{-1}C_{L}C_{\Theta}\log c_{\Theta}\right)^{2}\left(\tilde{\mathbb{P}}l_{\theta}^{(0)}\right)^{-2}\frac{(1+\zeta)^{2(T_{1}-1)-1}}{2\zeta+\zeta^{2}}\\ &\leq C_{\mathrm{exp}}\varepsilon^{-2}-c_{\mathrm{exp}}.\end{aligned}$
holds for some positive constants $C_{\mathrm{exp}}$ and $c_{\mathrm{exp}}$ that do not depend on $\varepsilon$. So we have proven the last inequality in Equation˜9.
Next, we consider the second inequality in Equation˜9. When $\pi^{\prime}$ is a constant policy, Theorem˜B.4 implies that for sufficiently small $\varepsilon$, $T^{*}(\pi^{\prime},\varepsilon)=\infty$ and the inequality holds vacuously. Otherwise, for a polynomial growth policy $\pi^{\prime}$ under Assumptions˜B.1, B.2 and B.3, there exist $\Theta$, $\Theta^{*}$, $l(x;\theta)$, and $\mathcal{R}(x)$ such that for $\theta^{(t)}_{\pi^{\prime}}$,
$\tilde{\mathbb{P}}l_{\theta^{(t+1)}_{\pi^{\prime}}}-\tilde{\mathbb{P}}l_{\theta^{(t)}_{\pi^{\prime}}}=\min\left\{-\frac{\gamma c}{4\beta}\tilde{\mathbb{P}}l_{\theta^{(t)}_{\pi^{\prime}}}+\rho\gamma\left(2-\frac{c}{4\beta}\right)2C_{L}C_{\Theta}\log c_{\Theta}\frac{1}{\sqrt{n_{t}^{\prime}}},0\right\},$
for some $0<\rho\leq 1$, and by Assumption˜B.3,
@@
Besides,
$\tilde{\mathbb{P}}l_{\theta^{(T^{*}(\pi^{\prime},\varepsilon))}_{\pi^{\prime}}}\leq c_{3}^{-1}\left(r^{*}-\mathbb{E}_{x\sim\mathcal{P}_{\theta^{(T^{*}(\pi^{\prime},\varepsilon))}_{\pi^{\prime}}}}[\mathcal{R}(x)]\right)\leq c_{3}^{-1}\varepsilon,$
@@
so the second inequality in Equation˜9 also holds.
@@
B.5 Extension to Rewards with Random Noise
Our theoretical conclusions directly generalize to cases with noisy rewards. Consider a perturbed reward function $\hat{\mathcal{R}}(x)=\mathcal{R}(x)+\varepsilon$, where $\varepsilon$ is zero-mean noise independent of $\mathcal{R}(x)$. Crucially, the expected reward $r(\theta)$ remains unchanged, preserving all key theoretical results. Although noisy rewards affect the sample selection process, the concentration bounds for selected sample sizes (Section˜B.4) and the expected reward under the selected distribution remain valid asymptotically. This robustness stems from the exponential policy’s inherent adaptivity by progressively increasing sample sizes (Theorem˜4.3), it naturally compensates for reward noise variance.
Appendix C Diffusion Probabilistic Models
To model the reverse process, the time interval $[0,1]$ is discretized into steps $0=\tau_{0}<\tau_{1}<\cdots<\tau_{K}=1$. Starting at $\tau_{K}$, the reverse diffusion process gradually removes noise step by step until reaching $\tau_{0}$. At each step $k$, the noise added during the forward process is estimated and removed using a trained noise prediction model $\bm{\varepsilon}_{\theta}(\cdot,\tau_{k})$. This model learns to predict the noise $\bm{\varepsilon}$ based on the noisy sample $\bm{x}_{\tau_{k}}$, enabling the reconstruction of cleaner samples at each step.
where $\mathscr{D}$ denotes the distribution of training data.
Appendix D Algorithm for Realistic Experiments
We follow the algorithm in Algorithm˜1. Empirically, in the two realistic settings (image denoising and math reasoning), our iterative learning policy satisfies that $n_{t}$ is an integer multiple of the batch size. Specifically,
@@ Appendix E Implementation Details
We provide the code for all the experiments in the GitHub repository: https://github.com/zylipku/spend-wisely.
@@ E.1 Toy Example
In this subsection, we extend the theoretical intuition from Theorem˜3.1 to high-dimensional settings and then derive the expected reward of each iteration, as well as the corresponding optimal expected reward.
@@
Under the MLE update, suppose that the generator is a $d$-dimensional Gaussian distribution $\mathcal{N}(\bm{\theta},\sigma^{2}\bm{I}_{d})$, the reward model is an exponential function $\mathcal{R}(\bm{x})=\exp\left(-\|\bm{x}\|^{2}/(2\kappa^{2})\right)$, and we start from
$\left\|\bm{\theta}^{(0)}\right\|\leq\left(1+\sigma^{2}/\kappa^{2}\right)^{T}\left(d\left(\sigma^{2}+\kappa^{2}\right)\right)^{1/2}.$
Considering the optimization problem
$\max_{\pi}$ $\quad\mathbb{E}_{\bm{\theta}^{(T)}}\left[\mathbb{E}_{\bm{x}\sim\mathcal{P}_{\bm{\theta}^{(T)}}}\mathcal{R}(\bm{x})\right]\quad\text{s.t.}\quad\sum_{k=0}^{T-1}n_{t}\leq C,$
@@
Since each component of $\bm{\theta}^{(t)}$ is i.i.d. and its distribution conditioned on the initial parameter $\theta^{(0)}$ and the policy $\pi$ is given by Lemma˜A.1, it follows that
$p\left(\bm{\theta}^{(T)}\mid\bm{\theta}^{(0)},\pi\right)=\mathcal{N}\left(\frac{\bm{\theta}^{(0)}}{(1+\sigma^{2}/\kappa^{2})^{T}},\sigma^{2}\sum_{t=0}^{T-1}\frac{1}{n_{t}(1+\sigma^{2}/\kappa^{2})^{2(T-t)-1}}\bm{I}_{d}\right)$
for a specific policy $\pi:\{n_{t}\}_{t}$. By the definition of the reward model, $\mathcal{R}(\bm{x})$ factors as
$\mathcal{R}(\bm{x})=\prod_{i=1}^{d}\mathcal{R}(x_{i}),\quad\bm{x}=(x_{1},x_{2},\cdots,x_{d}).$
Repeating the argument in the proof of Theorem˜3.1 shows that
@@
$=\left(\frac{\kappa}{\sqrt{\sigma^{2}+\kappa^{2}+\sigma_{T}^{2}}}\right)^{d}\exp\left(-\frac{\left\|\bm{\theta}^{(0)}\right\|^{2}}{2(\sigma^{2}+\kappa^{2}+\sigma_{T}^{2})(1+\sigma^{2}/\kappa^{2})^{2T}}\right).$
With assumption $\lVert\bm{\mu}_{T}\rVert\leq\left(d(\sigma^{2}+\kappa^{2})\right)^{1/2}$, the right side increases as $\left(\sigma^{2}+\kappa^{2}+\sigma_{T}^{2}\right)^{-1/2}$ increases. Hence, figuring with the proof of Theorem˜3.1 in Appendix˜A, one obtains
@@
Theorem˜E.1 therefore shows that the optimal iterative policy in the high-dimensional case mirrors the one-dimensional scenario. Next, we derive the expected reward in the high-dimensional case. Specifically, if $\bm{x}\sim\mathcal{N}(\bm{\theta},\sigma^{2}\bm{I}_{d})$, then
$\mathbb{E}_{\bm{x}\sim\mathcal{N}(\bm{\theta},\sigma^{2}\bm{I}_{d})}[\mathcal{R}(\bm{x})]$ $=\int\frac{1}{(2\pi\sigma^{2})^{d/2}}\exp\left(-\frac{\lVert\bm{x}\rVert^{2}}{2\kappa^{2}}-\frac{\lVert\bm{x}-\bm{\theta}\rVert^{2}}{2\sigma^{2}}\right)\mathrm{d}\bm{x}$
@@
E.2 Synthetic Experiment with Gaussian Distributions
Setup.
We begin with the synthetic setting introduced in our warm-up (Section˜3.2), which involves optimizing a generative distribution parameterized as a Gaussian to maximize the expected reward. Specifically, we focus on a two-dimensional space using MLE updates at each iteration. The generator is set to $\mathcal{N}(\bm{\theta},\bm{I}_{2})$, with fixed variance. The reward model is defined as an exponential function $\mathcal{R}(\bm{x})=\exp\left(-\lVert\bm{x}\rVert^{2}/(2\kappa^{2})\right)$, favoring data concentrated near the origin. $\kappa$ controls the flatness of the reward.
All theoretical results from the warm-up extend naturally to the high-dimensional case, as detailed in Section˜E.1. We now validate that with equal computational budgets and iterations, the exponential scheme outperforms the linear scheme, which outperforms the constant scheme.
Implementation.
We consider two settings with $\kappa^{2}\in\{2,4\}$. Based on theoretical analysis, the exponential scheme is set as $n_{t}=\lfloor n_{0}\cdot(1+1/\kappa^{2})^{t}\rfloor$, where $n_{0}=10$. For a fixed total budget and number of iterations, the constant scheme is then
$n_{t}=\left\lfloor T^{-1}\sum_{k=0}^{T-1}n_{0}\cdot(1+1/\kappa^{2})^{k}\right\rfloor,$
while the linear scheme is
$n_{t}=\left\lfloor\frac{2(t+1)}{T(T-1)}\sum_{k=0}^{T-1}n_{0}(1+1/\kappa^{2})^{k}\right\rfloor.$
The initial parameter is set to $\bm{\theta}^{(0)}=(1,1)^{\top}$, and the optimal solution for the reward function is $\bm{\theta}^{*}=(0,0)^{\top}$. The optimal reward value is then
$\mathbb{E}_{\bm{x}\sim\mathcal{N}(\bm{0},\bm{I})}[\mathcal{R}(\bm{x})]=\left(1+1/\kappa^{2}\right)^{-1}.$
We evaluate how fast the three schemes approach this optimal reward value.
@@ E.2 Synthetic Experiment with Gaussian Distributions
$L\left(\bm{\theta}^{(t)}\right)=|S|^{-1}\sum_{\bm{x}\in S}\left\|\bm{\varepsilon}_{\bm{\theta}^{(t)}}(\alpha_{\tau_{\bm{x}}}\bm{x}+\beta_{\tau_{\bm{x}}}\bm{\varepsilon}_{\bm{x}},\tau_{\bm{x}})-\bm{\varepsilon}_{\bm{x}}\right\|^{2}$
@@
E.3 Image Denoising
The pre-trained model we choose is a public checkpoint that can be accessed via https://huggingface.co/google/ddpm-cifar10-32. For all denoising experiments, we fix batch size $B=640$ and learning rate $5\times 10^{-5}$.
At each iteration, synthetic samples are generated on the fly, scored by a reward model
$\mathcal{R}(\hat{\bm{x}},\bm{x})=\frac{\text{PSNR}(\hat{\bm{x}},\bm{x})-r_{\min}}{r_{\max}-r_{\min}},$
and accumulated into a chunk until it reaches size $n_{t}$. The full training procedure is given in Algorithm˜2. Each configuration requires around 100 GPU-hours on an NVIDIA A800 cluster. Note that our setup is much more computationally intensive than the usual training of a generative model in that it takes $s$ times of network inferences to generate synthetic data for network training.
To compare increasing-schedule policies (particularly exponential growth), we report the computational costs in terms of floating-point operations (FLOPs), which scale linearly with the number of forward passes. For diffusion models, generating one sample uses $s$ denoising steps (i.e., $s$ forward passes), and each optimization step introduces an additional forward pass.
E.4 Math Reasoning
@@
6:   Sample a mini-batch $\mathcal{B}=\{(q_{j},a_{j})\}_{j}\subseteq\mathcal{D}$.
@@
$L(\theta)=\frac{1}{|\mathcal{B}|}\sum_{(\hat{a},q)\in\mathcal{B}}\log p(\hat{a}\mid q;f(\cdot,\theta))$
@@
We summarize the iterative math-reasoning algorithm in Algorithm˜3. Each configuration requires approximately 400 GPU-hours on an NVIDIA A800 cluster.
Hyperparameters are as follows: one epoch per iteration on the selected data; a constant learning rate of $10^{-7}$; batch size $B=256$; and roughly 1,0001{,}000 total generator update steps. During generation we use temperature $0.3$ for diversity, and temperature 0 at evaluation for accuracy. The maximum generation length is $512$. Prompts follow the few-shot setting of the GSM8k dataset in OpenCompass (Contributors, 2023). Our implementation builds on the public OpenRLHF framework (Hu et al., 2024).